OpenStudy (anonymous):

Able can harvest a strawberry crop in 4 days. Barney can do it in 5 days. Charlie would take 6 days. If they all work together, how long will it take them to complete the harvest?

OpenStudy (saifoo.khan):

\[\frac14 + \frac15 + \frac16 = \frac 1x\]Solve for x.

OpenStudy (saifoo.khan):

Where 1/4 is the able's part 1/5 is the barney's part 1/6is the charlie's part

OpenStudy (lgbasallote):

let x = able's time y = barney's time z = charlie's time 1/4 + 1/5 + 1/6 = 1/t t/4 + t/5 + t/6 = 1 15t/60 + 12t/60 + 10t/60 = 1 37t/60 = 1 37t = 60 t = 60/37

OpenStudy (saifoo.khan):

that's in days @asker.

OpenStudy (lgbasallote):

riiiiiight